A certain factory manufactures rectangular papers. Various sizes are produced by increasing the length by 6 centimeters each second and by decreasing its width by 5 centimeters each second. At a certain instant, the length of the paper is 3 centimeters, and the width is 8 centimeters. What is the rate of change of the area of the paper at that instant?
1
Expert's answer
2020-12-08T19:06:35-0500

Change in length per second is "\Delta L=+6cm" and Change in width is "\Delta W=-5cm".

At the given instant "(L,W)=(3,8)", hence area of paper is "A_1=24cm^2"and area at next second will be "A_2=(3+6)\times (8-5)=27cm^2" .

Change in area per second is "\Delta A=A_2-A_1=3cm^2"

Hence Rate of change of area is "3cm^2\/sec" .